-2 x - 2 y =5
8 x+ 8 y=-20
Having a hard time with these. thanks for all help!

Answer by stanbon(75887)

-2 x - 2 y =5
8 x+ 8 y=-20
---------------------
Multiply thru 1st by 4 to get:
-8x -8y = 20
-------------------
Add that to the 2nd equation to get
0 + 0 = 0
----------------
That means the two equations are really the same.
Solve one of them for y and you get y = -x-(5/2)
Solution: (x , -x-(5/2)) for x = any Real Number
=================================
Cheers,
Stan H.